OpenStudy (anonymous):

The description below represents Function A and the table represents Function B. Function A The function is 5 more than 3 times x. Function B x y -1 2 0 5 1 8 Which statement is correct about the slope and y-intercept of the two functions?

OpenStudy (anonymous):

Their slopes are equal but y-intercepts are not equal. Their slopes are not equal but y-intercepts are equal. Both slopes and y intercepts are equal. Neither slopes nor y intercepts are equal.

OpenStudy (anonymous):

i don't get your function b

OpenStudy (anonymous):

does all of those numbers belong to y?

OpenStudy (anonymous):

x y -1 2 0 5 -1 8

OpenStudy (anonymous):

function a y=3x+5 function b y=3x+5 (I'm assuming you meant 1,8 for that last coordinate) so both slopes and y-intercept is are equal do you know how to form an equation from those coordinates though?

OpenStudy (anonymous):

Yeah it was 1,8 I have no clue how to do any of this

OpenStudy (anonymous):

i used slope intercept formula to figure out the function b so y-intercept is (0,y) that 0 is always there whenever we're talking about y-intercept. we know from the function b that one coordinate is (0,5) so we plug that into the slope intercept form y=mx+5 we are also given another point (1,8) We can plug that into our equation 8=m(1)+5 then, solve for m 8-5=3 m=3 so the equation for function b is y=3x+5

OpenStudy (anonymous):

So the answer is C?

OpenStudy (dan815):

tounge licker

OpenStudy (anonymous):

Wtf?